# Affine Chabauty I

### v1

@@ 7.2. Hyperelliptic curves over imaginary quadratic fields
where $f(x)\in\mathcal{O}_{K}[x]$ is a squarefree polynomial of degree $2g+2\geq 4$ whose leading coefficient is a square in $\mathcal{O}_{K}$. Let $\mathcal{Y}\subseteq\mathbb{A}^{2}_{\mathcal{O}_{K}}$ be the model over $\mathcal{O}_{K}$ defined by the same equation. Write $Y=X\smallsetminus D$ and $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ with $X$ and $\mathcal{X}$ given by taking the closure in the weighted projective space $\mathbb{P}(1,g+1,1)$. Assume that $r=\operatorname{\mathrm{rk}}J(K)$ is equal to  $g$. Since the degree of $f$ is even and the leading coefficient is a square, there are two points at infinity, which are both $K$-rational. Thus (5.1) is satisfied for empty $S$ and we may apply ˜5.1 to determine the $\mathcal{O}_{K}$-points of $\mathcal{Y}$. Note that the model $\mathcal{X}/\mathcal{O}_{K}$ defined by the equation (7.2) might not be regular, in which case we have to pass to a desingularisation $\mathcal{X}^{\prime}\to\mathcal{X}$ in order to specify the reduction types as explained in Section˜3.4. Here we only use a single prime $\mathfrak{p}$; one could also use all primes lying over a fixed rational prime as in ˜5.4 and [GM25, §5].
@@
It is convenient to choose an auxiliary prime $\mathfrak{p}$ with $K_{\mathfrak{p}}=\mathbb{Q}_{p}$. We choose $\mathfrak{p}\coloneqq(2-\sqrt{-3})$ which divides the rational prime $7$. The curve has rank 1 over $\mathbb{Q}$ but rank $2$ over $K=\mathbb{Q}(\zeta_{3})$. For the lower bound $\operatorname{\mathrm{rk}}J(K)\geq 2$ it suffices to check that
@@
$\mathcal{Y}(\mathcal{O}_{K,\mathfrak{p}})\to K_{\mathfrak{p}},\quad P\mapsto\int_{P_{0}}^{P}\omega$
vanishes on $\mathcal{Y}(\mathcal{O}_{K})$. There is no constant in the Chabauty function since the base point $P_{0}=(0,1)$ is itself an $\mathcal{O}_{K}$-integral point. The annihilating differential $\omega$ is a nontrivial linear combination of the three basis differentials $\omega_{0},\omega_{1},\omega_{2}$ of $\mathrm{H}^{0}(X,\Omega^{1}(D))$, where $\omega_{j}=x^{j}\frac{\mathrm{d}x}{y}$. Using the known integral points $P_{0},P_{1},P_{2}$, we find that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{K,\mathfrak{p}})\to K_{\mathfrak{p}}$ defined by
@@
vanishes on $\mathcal{Y}(\mathcal{O}_{K})$. It does not necessarily vanish on $K$-rational points. For example for $P=(33/20,4073/8000)$, which is not integral at $2$ and $5$, one has $\rho(P)\neq 0$ in $\mathbb{Q}_{7}$. Expanding the determinant in (7.5) along the last row, we see that an annihilating log differential is given by $\omega=\alpha_{0}\omega_{0}+\alpha_{1}\omega_{1}+\alpha_{2}\omega_{2}$ with coefficients

### v2

@@ 7.3. Hyperelliptic curves over imaginary quadratic fields
where $f(x)\in\mathcal{O}_{K}[x]$ is a squarefree polynomial of degree $2g+2\geq 4$ whose leading coefficient is a square in $\mathcal{O}_{K}$. Let $\mathcal{Y}\subseteq\mathbb{A}^{2}_{\mathcal{O}_{K}}$ be the model over $\mathcal{O}_{K}$ defined by the same equation. Write $Y=X\smallsetminus D$ and $\mathcal{Y}=\mathcal{X}\smallsetminus\operatorname{\mathcal{D}}$ with $X$ and $\mathcal{X}$ given by taking the closure in the weighted projective space $\mathbb{P}(1,g+1,1)$. Assume that $r=\operatorname{\mathrm{rk}}J(K)$ is equal to  $g$. Since the degree of $f$ is even and the leading coefficient is a square, there are two points at infinity, which are both $K$-rational. Thus (5.1) is satisfied for empty $S$ and we may apply ˜5.1 to determine the $\mathcal{O}_{K}$-points of $\mathcal{Y}$. Note that the model $\mathcal{X}/\mathcal{O}_{K}$ defined by the equation (7.2) might not be regular, in which case we have to pass to a desingularisation $\mathcal{X}^{\prime}\to\mathcal{X}$ in order to specify the reduction types as explained in Section˜3.4. Here we only use a single prime $\mathfrak{p}$; one could also use all primes lying over a fixed rational prime as in ˜5.3 and [GM25, §5].
@@
It is convenient to choose an auxiliary prime $\mathfrak{p}$ with $K_{\mathfrak{p}}=\mathbb{Q}_{p}$. We choose $\mathfrak{p}\coloneqq(2-\sqrt{-3})$ which divides the rational prime $7$. The Jacobian has rank 1 over $\mathbb{Q}$ but rank $2$ over $K=\mathbb{Q}(\zeta_{3})$. For the lower bound $\operatorname{\mathrm{rk}}J(K)\geq 2$ it suffices to check that
@@
$\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}},\quad P\mapsto\int_{P_{0}}^{P}\omega$
vanishes on $\mathcal{Y}(\mathcal{O}_{K})$. There is no constant in the Chabauty function since the base point $P_{0}=(0,1)$ is itself an $\mathcal{O}_{K}$-integral point. The annihilating differential $\omega$ is a nontrivial linear combination of the three basis differentials $\omega_{0},\omega_{1},\omega_{2}$ of $\mathrm{H}^{0}(X,\Omega^{1}(D))$, where $\omega_{j}=x^{j}\frac{\mathrm{d}x}{y}$. Using the known integral points $P_{0},P_{1},P_{2}$, we find that the function $\rho\colon\mathcal{Y}(\mathcal{O}_{\mathfrak{p}})\to K_{\mathfrak{p}}$ defined by
@@
vanishes on $\mathcal{Y}(\mathcal{O}_{K})$. We remark that it does not necessarily vanish on $K$-rational points. For example for $P=(33/20,4073/8000)$, which is not integral at $2$ and $5$, one has $\rho(P)\neq 0$ in $\mathbb{Q}_{7}$. Expanding the determinant in (7.5) along the last row, we see that an annihilating log differential is given by $\omega=\alpha_{0}\omega_{0}+\alpha_{1}\omega_{1}+\alpha_{2}\omega_{2}$ with coefficients